Grade 11Physical Chemistry

The volume of carbon dioxide gas evolved at STP by heating 7.3 gm of Mg(HCO3)2 will be?

Ssssd
8 Years agoGrade 11

1 Answer

ADARSH SINGH
8 Years ago
 on heating magnisium hydrogen carbonate, 
 following reaction takes place-
 
                     Mg(HCO3)2 = MgO + H2O + 2CO2
 
NOW , TOTAL MOLECULAR MASS OF Mg(HCO3)2 IS 146 AND THAT OF C02 IS 44.
 
THEREFORE, NO OF MOLES OF Mg(HCO3) PRESENT IN OUR REACTION= 7.3/146 = 0.05
 
THEREFORE BY UNITARY METHOD,
 
NO OF MOLES FORMED OF CO  = 2 X 0.05 = 0.1 MOLE.
 
NOW, WE KNOW THAT IN STP, 1 MOLE OF ANY GAS HAS VOLUME 22.4L
 
THEREFORE VOLUME OF CO2 PRODUCED = 22.4 X 0.1 = 2.24 L
 
HOPE YOU GOT YOUR ANSWER .
IN CASE OF ANY DOUBT, PLEASE COMMENT BACK.
 
>>>>